OpenStudy (anonymous):

Which numbers are solutions of the equation?|1/2x|=9 A. {18, –18} B. {18, 18} C. {–18, –11} D. {18, –11}

OpenStudy (shamil98):

1/2x = 9 multiply both sides 2/1 or 2 x = 18 1/2x = -9 now do the same for this.

OpenStudy (anonymous):

so a?

OpenStudy (agent0smith):

Is it \[\Large \left| \frac{ 1 }{2x } \right|= 9\]or \[\Large \left| \frac{ 1 }{2 } x \right|= 9\]

OpenStudy (agent0smith):

Guessing the second one.

OpenStudy (shamil98):

yes. it is a.